The beam height in the cavity became totally different from the previous one and the shims needed to be much thicker than before. This is probably because of the alignment of the newly-glue curved mirror.
As the beam height is 2~2.5mm higher, two shims need to be stacked. The preliminary check of the heights using the alignment disks (dummy PDs) suggested the following combinations.
QPD1(SHORT) D1201467-03 (SN 007) + D1201467-03 (SN 008) (2.0 mm + 2.0 mm = 4 mm)
QPD2(LONG) D1201467-01 (SN 001) + D1201467-01 (SN 002) (1.5 mm + 1.5 mm = 3 mm)
DCPD1(TRANS) D1201467-02 (SN 006) + D1201467-03 (SN 005) (1.75mm + 2.0 mm = 3.75 mm)
DCPD2(REFL) D1201467-02 (SN 002) + D1201467-03 (SN 006) (1.75mm + 2.0 mm = 3.75 mm)
This resulted that the fixing button head socket screws for the PD housings to be replaced from 5/16" to 7/16". Stephen kept CLASS A spare screws from Jeff's time.
For the DCPD alignment, a cap-removed Excelitas 3mm InGaAs PD is used. -> This needs to be returned to the PD stock next time.
- DCPD1 was aligned using the zoomed CCD image (Attachment 1). Once the beam is aligned, the angle was tweaked to have the reflection nicely dumped by the glass beam dump (Attachment 2).
- DCPD2 was aligned too. (Attachment 2/3)
- The two housings were fastened by a torque wrench at 2 inch lb.
Continue with the QPDs. The QPD amp was already set.
The cable of the CCD monitor has a problem -> need to check what's wrong
The servo box probably have large offset at the output stage or somewhere (but not input stages).